The sum of the first 4 terms of an arithmetic sequence is 24. The 7th term exceeds the 4th term by 24. Determine the sum of the first 12 terms
Answer by alicealc(293)

The formula for the nth term in an arithmetic sequence is:
Un = U1 + (n - 1)d
in which Un is the nth term, U1 is the 1st term, and d is the difference.
So, for the 2nd term: U2 = U1 + (2 - 1)d = U1 + d
for the 3rd term: U3 = U1 + (3 - 1)d = U1 + 2d
for the 4th term: U4 = U1 + (4 - 1)d = U1 + 3d
and so on.

The sum of the first 4 terms of an arithmetic sequence is 24:
U1 + U2 + U3 + U4 = 24
U1 + (U1 + d) + (U1 + 2d) + (U1 + 3d) = 24
4U1 + 6d = 24
divide all by 2:
2U1 + 3d = 12

The 7th term exceeds the 4th term by 24:
7th term: U7 = U1 + (7 - 1)d = U1 + 6d
U7 = U4 + 24
U1 + 6d = U1 + 3d + 24
U1 + 6d - U1 - 3d = 24
6d = 3d + 24
6d - 3d = 24
3d = 24
d = 24/3 = 8

2U1 + 3d = 12
2U1 + 3(8) = 12
2U1 + 24 = 12
2U1 = -12
U1 = 12/2 = -6

The sum of the first n term can be found by using this formula:
Sn = n/2 * {2U1 + (n - 1)d}

so, the sum of the first 12 term is:
S12 = 12/2 * {2(-6) + (12 - 1)8}
S12 = 6 * (-12 + 11*8)
S12 = 6 * (-12 + 88)
S12 = 6 * 76
S12 = 456